OpenStudy (anonymous):

A rectangular field bounded on Side by a lake is to be fenced on 3 sides by 800 m of fence. What demensions wil produce a maximum area?

OpenStudy (amistre64):

what formulas or functions would you use to develop relationships between side lengths?

OpenStudy (amistre64):

 2x + y = 800 xy = Area to maximize ...

OpenStudy (amistre64):

using calculus: A' = xy' + x'y if we let A wrt/x, then x'=1 A' = xy' + y, this is min or max when A=0 0 = xy' + y -y/x = y' from P we get 2+y' = 0, so when y'=-2 -y/x = -2 ... or when y=2x 2x + 2x = 800; x = 800/4, y = 800/2